Fortes Inc. has provided the following data concerning one of the products in its standard cost system. Variable manufacturing overhead is applied to products on the basis of direct labor-hours. Standard Quantity or Hours per Unit of Output 8.7 ounces 0.6 hours 0.6 hours Standard Price or Rate Inputs Direct materials Direct labor Variable manufacturing overhead 6.80 per ounce $28.70 per hour s 5.10 per hour 48 The company has reported the following actual results for the product for April: 6,300 units S5,140 ounees Actual output Raw materials purchased Actual cost of rav materials purchased Raw materiala used in production Actual direct labor-hours Actual direct labor eost Actual variable overhead cont 316,540 4,820 ounces 3,530 hours $105,850 17,090 Required: a. Compute the materials price variance for April b. Compute the materials quantity variance for April. c. Compute the labor rate variance for April d. Compute the labor efficiency variance for April. e. Compute the variable overhead rate variance for April. f. Compute the variable overhead efficiency variance for April. (Indicate the effect of each variance by selecting F for favorable, U for unfavorable, and None for no effect (i.e., zero variance). Input all amounts as positive values.) a. Materials price variance b. Materials quantity variance c. Labor rate variance d Labor efficienoy variance e. Variable overhead rate variance Variable overhead efficiency variance
Answer #1

Revised standard budget based on actual output

Particulars qty/hours rate Amount
Direct material (6300*8.7) 54810 ounces 6.80 372708
Direct labour (6300*0.6) 3780 hours 28.70 108486
Variable OH 3780 hours 5.10 19278

Actual expenses

Particulars qty/hours rate Amount
Direct material 54820 ounces 5.7406 314702
Direct labour 3530 hours 29.985 105850
Variable OH 3530 hours 4.84 17090

Rate of material purchased = 316540/55140 = 5.7406

a) material price variance = (std price - actual price ) * actual unit

= (6.80 - 5.74) * 54820 = 58109 favourable

b) material qty var = (std qty - actual qty) * std rate

= (54810-54820) * 6.8 = 68 adverse

c) labour rate variance = (std rate - actual rate ) * actual hours

= (28.70 - 29.985)*3530 = 4536 adverse

d) labour efficiency variance = (std hours - actual hourse) *std rate

= (3780 - 3530) * 2870 = 717500

e) variable ohrate varaince - =(5.10-4.84)*3530 = 917.8

f) VOEV = (3780 - 3530) * 5.10 = 1275